# Neural Fields for Adaptive Photoacoustic Computed Tomography

### v2

@@ 3.2 Leaf Phantom
We use the phantom data from [20], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in lower left). The phantom is surrounded by water with a temperature of 26∘C and the wavelength of laser illumination is 700 nm in the experiment. Similarly, the constant background SOS is set to 1499.4 m/s according to [37]. Due to nonideal transducer alignment and transducer impulse responses, the collected PA signals are calibrated using measured ring error and transducer electrical impulse response (EIR) before feeding into the algorithms. The same calibration is also applied to the following in vivo experiment.
The reconstruction results of NF-APACT and several other methods are shown in Fig. 5a. The lower left of the phantom is dim due to the nonuniform illumination [20]. We searched over SOS values to report the best outputs, using SOS values of 1505.0 m/s for conventional DAS and 1520.0 m/s for body SOS in Dual-SOS DAS. NF-APACT is able to refocus the leaf veins in the cylinder (blue) as well as inside the hole (red), where the signals are weak due to nonuniform and acoustic reflections, and DAS and APACT both fail. The SOS reconstructed by NF-APACT has a lower SOS region in the middle matching the hole filled with water, while APACT smooths out the SOS and cannot resolve the hole. In terms of compute time, our method is also an order of magnitude faster than APACT.
@@ 3.3 In vivo Mouse Liver
In our in vivo experiment, we use the PA signals collected from a nude mouse liver at 1064 nm in [19]. An illustration of the imaged section is shown in the lower left corner. The water temperature was 31∘C during the experiment, thus we used a constant background SOS of 1511.4 m/s. We used 1515.0 m/s for conventional DAS and 1560.0 m/s for Dual-SOS DAS.
The reconstructed results are shown in Fig. 5b. It is hard to fully characterize the reconstruction qualities without ground truth, but we are able to demonstrate the success of our method by looking at several important features. NF-APACT is able to increase the overall contrast and recover structures like blood vessels (green and blue) and bright bone structures (red). Our reconstructed SOS has a superior match to the liver’s anatomical shape. In comparison, APACT reconstructs a smooth SOS in which the middle, high-SOS region does not align well with the liver structure.
@@ 4 Discussion and Conclusion
We have introduced NF-APACT, a self-supervised framework achieving fast, state-of-the-art performance in joint reconstruction of initial pressure and SOS in PACT. This method combines the neural field representation from computer graphics with an extension of the optical metaphors of autofocus and adaptive optics used in previous methods to a multichannel image representation analogous to a focal stack. We highlight the following as the key features of our novel method:
@@
Reliability and interpretability. NF-APACT combines signal processing with neural networks to provide reasonable guess of the SOS without a black box. In contrast to deep learning methods where neural networks are trained on large datasets, our method utilizes a small MLP as an efficient representation in a physics-based framework. Our approach enhances reliability and interpretability, making it well-suited for scientific and medical applications where black-box models are often undesirable.
@@
Time-performance trade-offs. Existing reconstruction methods offer diverse solutions for different demands. Based on our benchmarking, we can recommend Dual-SOS DAS with a properly-selected SOS for real-time removal of many aberrations. When image quality is key, we offer a new method to access state-of-the-art reconstruction of initial pressure and SOS in roughly 1 minute. The number of training epochs and learning rate in NF-APACT further allow for customization of the time-performance trade-off. Additionally, we demonstrate 6-second reconstruction when SOS is measured directly.
@@
Generalizability. Although only demonstrated in 2D, our framework could potentially extend to 3D PACT systems. In 3D PACT, where 2D scans from various locations and orientations are combined to reconstruct a 3D volume, the major challenge lies in the large number of unknowns in the 3D SOS, which scales up the computational cost by order of magnitudes. To address this challenge, the efficient SOS representation combined with the efficient and differentiable forward model in NF-APACT offers a promising solution. By reducing the degree of freedoms and lowering computational costs, NF-APACT could enable practical 3D PACT applications.
Acknowledgment
We would like to thank Liujie Gu and Yan Luo for helping us with the numerical simulations.

### v3

@@ 4.2 Leaf Phantom
We use the phantom data from [6], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in lower left of Fig. 5a and details in Sec. S1.2). The reconstructions of our methods and the baselines are shown in Fig. 5a. The lower left of the phantom is dim due to nonuniform illumination [6]. We searched over SOS values to report the best outputs, using SOS values of 1505.0 m/s for conventional DAS and 1520.0 m/s for body SOS in Dual-SOS DAS. Ours are able to refocus the leaf veins in the cylinder (blue) as well as inside the hole (red), where the signals are weak and DAS and APACT both fail. The SOS reconstructed by our methods have a lower SOS region in the middle matching the hole filled with water, while APACT smooths out the SOS and cannot resolve the hole. PG still suffers from stripe-like artifacts in the SOS (zoom in) but also reconstructs a well-resolved IP image.
@@ 4.3 In vivo Mouse Liver
In our in vivo experiment, we used a nude mouse liver from [3] (details in Sec. S1.3). The reconstructions are shown in Fig. 5b along with a cartoon illustration of the imaged section (bottom left corner). We hand-tuned SOS assumptions to 1517.0 m/s for conventional DAS and 1560.0 m/s for Dual-SOS DAS. Despite lacking ground truth, we observe success on several important features. Our method with NF is able to increase the overall contrast and recover structures like blood vessels (green and blue) and bright spine and rib structures (red) in the IP image. Our reconstructed SOS has a superior match to the liver’s anatomical shape (see overlaid figure in Fig. S1). In comparison, APACT reconstructs a smooth SOS in which the middle, high-SOS region does not align well with the liver. Ours with PG also suffers from severe artifacts in the SOS.
4.4 Ablation Studies
Our ablation studies focus on the effects of the number of delays , TV regularization , and network size. We observe an increase in the performance at the cost of compute time as more extra delay distances are used and more redundancy is created in the DAS stack. This improvement gradually diminishes beyond a delay count of 16, which we selected for our implementation (see Fig. 6). For NF network size (see Fig. S3 and Sec. S3.3), our method fails either when the network is too small (NF’s implicit smoothness regularization is too strong) or large (the NF has a stronger ability to learn high-frequency features and has weaker implicit regularization). For the TV regularization of PG (see Fig. S2 and Sec. S3.2), a larger encourages smoother SOS results and reaches optimal performance at a weight of .
@@ 5 Discussion and Conclusion
We have introduced a self-supervised framework achieving fast, state-of-the-art performance in joint reconstruction of initial pressure and speed of sound in PACT. This method combines the neural field representation from computer graphics with an extension of the optical metaphors of autofocus and adaptive optics used in previous methods to a multichannel image representation analogous to a focal stack. We highlight the following as the key features of our novel method:
@@
Reliability and interpretability. Our method combines signal processing with neural networks to provide reasonable guess of the SOS without a black box. In contrast to deep learning methods where neural networks are trained on large datasets, our method utilizes a small MLP as an efficient representation in a physics-based framework and incorporates redundant information in the measurements to efficiently solve a partially-blind inverse problem. Our approach enhances reliability and interpretability, making it well-suited for scientific and medical applications where black-box models are often undesirable.
@@
Time-performance trade-offs. Existing reconstruction methods offer diverse solutions for different demands. Based on our benchmarking, we can recommend Dual-SOS DAS with a properly-selected SOS for real-time removal of many aberrations. When image quality is key, we offer a new method to access state-of-the-art reconstruction of the IP and SOS in roughly 40 seconds. The delay count, network size, and number of training epochs in NF further allow for customization of the time-performance trade-off. Additionally, we demonstrate 3-second reconstruction when SOS is measured directly.
@@
Generalizability. Although only demonstrated in 2D, our framework could potentially extend to 3D PACT systems [4]. In 3D PACT, where multiple 2D scans are combined to reconstruct a 3D volume, the major challenge lies in the large number of unknowns in the 3D SOS, which drastically scales up the computational time and memory requirements. To address this challenge, our efficient NF SOS representation combined with efficient and differentiable forward model offers a promising solution. By reducing the degrees of freedom and lowering computational costs, the proposed method could enable practical 3D PACT applications.
Supplementary Material
S1 Experiment Details
S1.1 Numerical Phantoms
In this section, we describe the simulation details of our numerical phantoms.
We use the same IP image shown in Fig. 4a for all numerical phantoms but use five SOS with different variations (see Fig. S4a). SOS values range from 1490.0 to 1650.0 m/s in our numerical dataset. The structures in the IP and SOS images are inspired by the mouse body111https://www.imaios.com/en/vet-anatomy/mouse/mouse-whole-body while the SOS values are based on realistic measurements222https://itis.swiss/virtual-population/tissue-properties/database/acoustic-properties/. For example, the small circles with high SOS in samples 1 and 3 correspond to bones, the orange circles in samples 1 and 2 correspond to the liver, and the circle with small SOS in samples 1, 4, and 5 correspond to the stomach lumen.
The background water SOS is set to 1499.4 m/s (assuming a 26∘C water temperature based on [20]) outside the sample. The PA signals are obtained by 2D numerical simulation using the k-Wave toolbox [32] assuming an evenly distributed 512-transducer ring array with a diameter of 10 cm (same setting for the leaf phantom and in vivo mouse liver experiments). Similar to [6], we take the derivative of the PA signals to simulate transducer effects: . The images are cropped into 75% overlapping 3.2mm 3.2mm patches, as for all reported results.
We show the IP and SOS reconstructions on all five numerical phantoms using our methods and APACT in Fig. S4 b,c,d. Results are qualitatively similar to the example discussed in the text, and numbers reported in tables are averaged over these five phantoms.
S1.2 Leaf Phantom
We use the experimentally collected phantom data from [6], which features a leaf in a 2.5 cm diameter agarose cylinder with a water-filled hole in the middle for acoustic heterogeneity (see photograph in Fig. 5a). The leaf phantom is surrounded by water with a temperature of 26∘C, and the wavelength of laser illumination is 700 nm in the experiment. The constant background SOS is set to 1499.4 m/s according to [20].
Due to nonideal transducer alignment and transducer impulse responses, the collected PA signals are calibrated using measured ring error and transducer electrical impulse response (EIR) before feeding into the algorithms. The same calibration is applied to the in vivo experiment.
S1.3 In vivo Mouse Liver
The PA signals in our in vivo experiments were collected from a nude mouse liver at 1064 nm [3]. The water temperature was 31∘C during the experiment, thus we used a background SOS of 1511.4 m/s. See figure S1 to see the alignment between recovered SOS and IP, which matches better in our method than others.
S1.4 Image Stitching
We adopted the same strategy used for APACT [6] in IP reconstruction. This process differs from most existing image reconstruction frameworks because the image is reconstructed patch-by-patch and stitched to form the final result. Images are cropped using 3.2 mm 3.2 mm Gaussian windows with an FWHM of 1.5 mm before being fed to our algorithm. The patch centers are arranged on a Cartesian grid, and adjacent patches overlap by 75%. With our proposed method for solving the speed of sound distribution by utilizing information from all patches, we obtain the wavefront correction model for each isoplanatic patch. This simultaneously yields the deconvolved, aberration-corrected image patches. These patches are then translated to their corresponding center positions and summed. The 75% overlap ensures a smooth transition during image stitching. Finally, the result is adjusted to remove the modulation pattern due to the Gaussian windows by dividing by the sum of the Gaussian window weights.
S2 Difference Between Our Wavefront Model and APACT’s Wavefront Model
One critical contribution of our work is the improved wavefront model compared to the previous work APACT [6]. As explained in the main text, the wavefront model represents the wavefront advancement or retardation induced by a nonuniform speed‐of‐sound distribution in tissue relative to an ideal spherical wave in a homogeneous medium. Mathematically, the wavefront function , which depicts the propagation length difference relative to the spherical wave when received by the transducers, is a function of the propagation direction . This wavefront function can be expanded using a Fourier series:
In APACT, only three low order harmonics were used, namely DC (constant ), , and . This is because, firstly, the first order terms and result in a shifted PSF, making them non-identifiable from the reconstructed images when the true IP is unknown. Secondly, the authors omitted the higher order terms, i.e. and for , to avoid the exponential increase in computation time during the exhaustive search for the optimal wavefront. However, as shown in Fig. 2, the actual wavefront function can be complex, which necessitate those high order terms. In our work, the wavefront functions corresponding to different isoplanatic patches are obtained from the learned SOS map. There is no explicit restriction on their profiles such as the assumption in APACT that they are slowly varying.
S3 Ablation Studies
We conducted ablation studies using our synthetic numerical phantoms on an NVIDIA RTX A6000 GPU. All results are averaged over 5 numerical phantoms, which share the same IP but have different SOS.
S3.1 Number of Delays
We show the performance of our method (with and without SOS oracle), pixel grid (with and without TV regularization), and APACT with respect to numbers of delays used in Fig. 6. We keep the range of the delay distances fixed at [-0.8 mm, 0.8 mm] and uniformly sample delay distances within this range based on the number of delays used. We show that more delay channels leads to increased performance and computation time for all methods. This originates from Eq. 7, where more delay channels promote the performance of pseudo-inverse deconvolution, provide more redundant information for SOS estimation, and require more computation time. Removing TV regularization in pixel grid (dashed green line) has little impact on the quality of the IP image but significantly decreases the quality of the SOS. Considering the performance and computation time, we choose 16 delays (marked with stars). We also show that the efficient parameterization of SOS gives NF a slight time advantage compared to PG (see lower right subplot), which takes more iterations to reach convergence.
S3.2 TV regularization for PG
In Fig. S2, we show the effects of TV regularization weight with our pixel grid SOS representation. We observe that the IP image quality drops when , while the SOS accuracy reaches optimum around . Therefore, we select (marked with a green star in Fig. S2) for all reconstructions in our implementation of PG.
S3.3 Network Size of NF
We compared the performance of our method with NF with different network sizes ranging from 131 to 264,707 parameters, shown in Fig. S3. We show that the implicit smoothness regularization of the MLP decreases as the network becomes larger, resulting in noisier SOS reconstructions and hence worse IP images. This can be observed from the comparison between no hidden layer (solid line) and one hidden layer (dashed line) and the overall decreasing trend of the two curves. An increase in network size also leads to a small rise in computation time. We select 0 hidden layer and 256 hidden features after an overall consideration of the performance and computation time (marked with a purple star).
S4 Supplementary Videos
We provide six videos showing the convergence process of the IP image and SOS for our methods with our numerical phantom (nf_numerical_phantom.mp4, and pg_numerical_phantom.mp4), leaf phantom (nf_leaf_phantom.mp4, and pg_leaf_phantom.mp4), and in vivo mouse liver (nf_in_vivo_mouse_liver.mp4 and pg_in_vivo_mouse_liver.mp4).